Find the slope and the y-intercept of the equation 2x − 3y = 6
Answer by Alan3354(69443)

Find the slope and the y-intercept of the equation 2x - 3y = 6
----------
Convert it to slope-intercept form, y = mx + b
m = slope, b = y-intercept.
------------
That means solve for y.
2x - 3y = 6
3y = 2x - 6
y = (2/3)x - 2
----
m = 2/3
b = -2